A force "A" is addes two force B =(i-2j+4k) and C=(2i+3j-6k) and given a unit vector force along the x axis as their resultant. Find the acceleration of the body 5 kg body , if A acts on the body


1
Expert's answer
2021-01-29T05:19:53-0500

"\overrightarrow{A}=\overrightarrow{B}+\overrightarrow{C}=(\overrightarrow{i}-2\overrightarrow{j}+4\overrightarrow{k})+(2\overrightarrow{i}+3\overrightarrow{j}-6\overrightarrow{k})="


"=3\overrightarrow{i}+\overrightarrow{j}-2\overrightarrow{k}"


"\overrightarrow{a}=\frac{\overrightarrow{A}}{m}=\frac{3\overrightarrow{i}+\overrightarrow{j}-2\overrightarrow{k}}{5}=0.6\overrightarrow{i}+0.2\overrightarrow{j}-0.4\overrightarrow{k}"


"|\overrightarrow{a}|=\sqrt{0.6^2+0.2^2+0.4^2}=0.75\ (m/s^2)"